# Power Diagram Enhanced Adaptive Isosurface Extraction from Signed Distance Fields

### v1

@@ 2 Related Work
2.1 Extraction on Pre-defined Grids
@@
2.2 Progressive Evolution Methods
@@ 3 Overview
4 Power Diagram-based Isosurface Extraction
@@ 4.1 Theoretical Foundation
4.2 Surface-Tangent Points Integration
@@
4.3 Stable Surface Extraction
@@
4.4 Localized Incremental Update
@@ 5.1 Pipeline
Step 1. Construct regular delaunay $\mathcal{D}$ from uniform grid vertices and their projections, obtain initial surface $\mathcal{M}$  (Section 4).
Step 2. Calculate the deviation metric $\delta$ between each triangle in $\mathcal{M}$ and the underlying surface (Section 5.2) and organize them in max-heap $\mathcal{Q}$.
Step 3. Extract triangle with maximum deviation $\delta_{c}$ from $\mathcal{Q}$, add corresponding new site $p_{c}$ and its projection $q_{c}$ (Section 5.3). Update $\mathcal{D}$ and $\mathcal{M}$ (Section 4.4), compute quality metrics for affected and newly added faces (Section 5.2) and insert them into $\mathcal{Q}$.
@@
5.2 Isosurface Approximation Quality Evaluation
@@
5.3 Strategic Site Insertion for Local Refinement
@@
Figure 7: Comparison with state-of-the-art methods on three high-genus inputs. Previous methods and Ours1 use uniform $50^{3}$ grid samples, while Ours2 starts from an $8^{3}$ grid and applies our adaptive sampling strategy until reaching 62500 samples. Our method demonstrates better topological preservation, with further improvements achieved through our adaptive sampling strategy.
6 Evaluation
6.1 Experiment Setting
@@
Our experimental evaluation was conducted on a system equipped with an AMD Ryzen 9 5950X CPU and 72 GB RAM. The test dataset comprised 100 models curated from Thingi10k and models previously used in [26] and [27], supplemented with several custom-designed models. All models were normalized to fit within a unit cube of $[-1/2,1/2]^{3}$, while the surface extraction was performed in the extended domain of $[-1,1]^{3}$.
Comparison Methods
@@
This study presents a comparative analysis between our approach and five state-of-the-art methods: three recent techniques—RFTA (REACH for Arc) [26], RFTS (REACH for Sphere) [27], and MCGrids [24]—and two classical methods—Marching Cubes [16] and NDCx (Neural Dual Contouring) [3].
@@
RFTA and RFTS utilize optimization-based approaches that exploit the tangential relationship between sample-centered spheres and surfaces for geometric reconstruction. Here, each sphere is centered at a sample point with its radius defined by the distance to the surface. For RFTS, we employed the default parameters provided in the authors’ GPyToolbox implementation [25], while RFTA was configured with 20 fine-tuning iterations and 30 maximum points per sphere. MCGrids employs an iterative isosurface extraction approach, whereas Marching Cubes and NDCx implement uniform spatial discretization strategies for surface extraction.
@@
We categorize existing methods into two distinct classes: Direct Extraction Methods, which utilize fixed information for isosurface extraction without additional input during reconstruction (including Marching Cubes, NDCx, RFTS, and RFTA), and Iterative Refinement Methods, which progressively insert new points and adjust results during the extraction process (McGrids).
6.2 Comparison with Direct Extraction Methods
@@
Figure 6 illustrates comparative analyses at uniform grid resolutions of $30^{3}$ and $50^{3}$ over the $[-1,1]^{3}$ domain, where all algorithms receive identical grid point SDF values and gradient information for low-genus model reconstruction. Traditional methods show distinct limitations: Marching Cubes and Neural Dual Contouring utilize only SDF sign information, resulting in poor detail preservation; RFTS cannot handle topological variations, restricting its application to simple models; and RFTA, despite improving upon its predecessor, fails to recover intricate surface details. Our method leverages projection points derived from gradient information with regular Delaunay-based extraction to achieve superior reconstruction quality. This is demonstrated in Figure 6 with the strawberry model and in Figure 2 with the golf ball model, where our approach successfully reconstructs surface cavities that RFTA fails to capture. Additionally, at $50^{3}$ resolution, RFTA’s tower reconstruction exhibits notable spherical artifacts.
@@
For high-genus models (Figure 7), we present two implementations: Our1, which uses standard $50^{3}$ grid inputs, and Our2, which employs an adaptive approach starting from an $8^{3}$ grid and expanding to $50^{3}$ total points (including both sample points and their projections) using the strategy detailed in Section 5. While conventional methods (Marching Cubes, NDCx, RFTS) fail to preserve intricate details such as chair backs, and RFTA shows only modest improvement, both of our implementations demonstrate better detail preservation. Our2, through its adaptive point sampling strategy, achieves particularly significant improvements in detail preservation.
@@
6.3 Comparison with Progressive Refinement Methods
@@
Additionally, we evaluated the reconstruction accuracy by measuring the L2 Chamfer distance between the extracted isosurfaces and ground truth across our dataset of 100 models. We employed two strategies: Ours1 used uniform $50^{3}$ grid samples as input, while Ours2 utilized our incremental construction approach until reaching 50,000 sampling points. Classical methods used uniform $50^{3}$ grid information as input, while MCGrids progressively added points until reaching 50,000 samples. The resulting average L2 Chamfer distances were: Marching Cubes: 2.143×10-4, RFTA: 0.421×10-4, Ours1: 0.209×10-4, McGrids: 9.49×10-6, and Ours2: 8.47×10-6. These quantitative results further validate our method’s superior isosurface extraction capabilities, demonstrating improved accuracy in both uniform sampling and adaptive refinement scenarios.
6.4 Other Experiments
For all results presented in this paper, the running time consistently remains around 10 seconds. Additional timing data, ablation studies, and visualizations of intermediate stages in the isosurface extraction process can be found in the Supplementary Material.
7 Conclusion, Limitations and Future Work
@@
Appendix A Iterative Evolution Analysis
Figure 13 demonstrates the iterative progression of our algorithm on two representative examples: a high-genus model and a model with intricate surface details. Starting from a uniform initialization with $15^{3}$ sampling points, the algorithm iteratively incorporates new sites until reaching 50,000 sites. For each intermediate stage, the highlighted sites indicate the newly added positions during the transition to the subsequent stage. The visualization reveals two primary functions of our site addition strategy: (1) Topology correction - as evidenced in the first model, where new sites are strategically placed to either establish necessary connections or separate incorrectly merged regions; (2) Geometric refinement - as demonstrated in the second model, where the progressive addition of sites continuously enriches surface details. These results validate the effectiveness of our iterative strategy in progressively improving both topological accuracy and geometric fidelity.
Appendix B Ablation Study
To evaluate the key components of our method, we conducted ablation studies on two critical aspects: 1) the effectiveness of incorporating projection points in the isosurface extraction process, and 2) the impact of our quality-driven iterative point addition strategy.
[…]
Figure 9: Comparison of results using (a) sampling points only versus (b) our complete method with both sampling points and their surface projections. Starting from a uniform 20³ grid and refining to 40,000 points, the projection-enhanced approach demonstrates superior topological preservation.
B.1 Surface Projection Integration
The addition of projection points enhances the algorithm’s ability to ”sense” the surface more intuitively during isosurface extraction, which is crucial when handling complex topologies and rich geometric details. Figure 9 demonstrates the results using a uniform $15^{3}$ grid as input, iteratively refined to 40,000 sampling points with: (a) our method using only sampling points without their projections, and (b) our complete method utilizing both sampling points and their surface projections.
The comparison clearly shows that the projection-enhanced approach preserves more complete topological structures, while the baseline method without projections exhibits numerous topological inaccuracies. This demonstrates how surface projections provide critical geometric guidance that significantly improves the algorithm’s ability to reconstruct complex features, particularly in regions with intricate topology or thin structures that might otherwise be missed when relying solely on discrete sampling points.
[…]
B.2 Quality-Driven Adaptive Refinement
To evaluate our iterative point addition strategy, we conducted experiments using different initial uniform grid resolutions while maintaining a fixed total of 40,000 sampling points. Figure 10 demonstrates that when a greater proportion of points are determined by our quality-driven strategy rather than uniform initialization, the reconstruction exhibits significantly better topological preservation. In contrast, models reconstructed primarily with uniformly distributed points show compromised topology, confirming that our adaptive approach more effectively allocates sampling points to topologically and geometrically critical regions.
[…]
Appendix C Computational Efficiency Analysis
Figure 11 illustrates the algorithm’s running time with varying numbers of iteratively added points using a complex model as an example. Due to our algorithm’s localized update approach, the time required to insert a new point remains relatively constant regardless of the total number of points already present in the regular Delaunay tetrahedralization. This locality property ensures that our method’s computational efficiency scales approximately linearly with increasing point count, rather than exhibiting higher complexity. The results validate the effectiveness of our localized update strategy, confirming the computational efficiency of our incremental framework.
[…]
Figure 12: Comparison of isosurface extraction results before (left) and after (right) post-processing. The application of Laplacian smoothing followed by gradient-based reprojection significantly enhances surface smoothness and overall quality while preserving the captured topology.
Appendix D Post-Processing Enhancements
In previous experiments, we validated our algorithm’s effectiveness in reconstructing complex topological structures. We further discovered that the surface quality can be significantly improved through a simple post-processing step. When we apply Laplacian smoothing to our reconstruction results (where topology is already largely correct) and then reproject all surface vertices back onto the zero-level set using gradient information, the overall quality of the reconstructed surface is further enhanced.
As shown in Figure 12, the surface after smoothing and reprojection appears noticeably smoother than the original result while maintaining topological accuracy. It is important to note that for fair comparison with other methods, we did not apply this post-processing step to any results presented in the main body of the paper.
[…]

### v2

@@ Abstract
Index Terms:
Signed Distance Fields, Isosurface Extraction, Delaunay Tetrahedralization, Surface Reconstruction.
@@ II Related Work
II-A Extraction on Pre-defined Grids
@@
II-B Progressive Evolution Methods
@@ III Overview
IV Power Diagram-based Isosurface Extraction
@@ IV-A Theoretical Foundation
IV-B Surface-Tangent Points Integration
@@
IV-C Stable Surface Extraction
@@
IV-D Localized Incremental Update
@@ V-A Pipeline
Step 1. Construct regular delaunay $\mathcal{D}$ from uniform grid vertices and their projections, obtain initial surface $\mathcal{M}$  (Section IV).
Step 2. Calculate the deviation metric $\delta$ between each triangle in $\mathcal{M}$ and the underlying surface (Section V-B) and organize them in max-heap $\mathcal{Q}$.
Step 3. Extract triangle with maximum deviation $\delta_{c}$ from $\mathcal{Q}$, add corresponding new site $p_{c}$ and its projection $q_{c}$ (Section V-C). Update $\mathcal{D}$ and $\mathcal{M}$ (Section IV-D), compute quality metrics for affected and newly added faces (Section V-B) and insert them into $\mathcal{Q}$.
@@
V-B Isosurface Approximation Quality Evaluation
@@
It is worth noting that this gradient-alignment metric might theoretically yield a favorable quality score for a normal-aligned triangle that is positioned at a distance from the actual zero-level set. However, our algorithm incorporates two complementary mechanisms that prevent such cases: (1) Our surface projection step consistently maintains triangles in close proximity to the zero-level set through the use of surface-tangent points, and (2) In the unlikely event that such a misplaced triangle exists, adjacent triangles would exhibit high energy values, which would automatically trigger point insertion during the refinement process to correct the local geometry.
V-C Strategic Site Insertion for Local Refinement
@@
Figure 7: Comparison with state-of-the-art methods on two high-genus inputs. Previous methods and Ours1 use uniform $50^{3}$ grid samples, while Ours2 starts from an $8^{3}$ grid and applies our adaptive sampling strategy until reaching 62500 samples. Our method demonstrates better topological preservation, with further improvements achieved through our adaptive sampling strategy.
VI Evaluation
VI-A Experiment Setting
@@
Our experimental evaluation was conducted on a system equipped with an AMD Ryzen 9 5950X CPU and 72 GB RAM. The test dataset comprised 100 models curated from Thingi10k and models previously used in [19] and [18], supplemented with some custom-designed models. Following RFTS’s [18] experimental design, all models were normalized to fit within a unit cube of $[-1/2,1/2]^{3}$, while the surface extraction was performed in the extended domain of $[-1,1]^{3}$. Our comparative experiments focused on two principal scenarios: (1) extraction under fixed input conditions (the core application scenario of RFTS/RFTA) to validate extraction capability, and (2) extraction in continuous fields.
@@
Comparison Methods. Our experimental evaluation benchmarks the proposed approach against five representative isosurface extraction techniques from the literature: three recent techniques—RFTA (REACH for Arc) [19], RFTS (REACH for Sphere) [18], and MCGrids [20]—and two classical methods—Marching Cubes [30] and Dual Contouring [14].
@@
RFTA and RFTS utilize optimization-based approaches that exploit the tangential relationship between sample-centered spheres and surfaces for geometric reconstruction. Here, each sphere is centered at a sample point with its radius defined by the distance to the surface. For RFTS, we employed the default parameters provided in the authors’ GPyToolbox implementation [31], while RFTA was configured with 20 fine-tuning iterations and 30 maximum points per sphere. MCGrids employs an iterative isosurface extraction approach, whereas Marching Cubes and Dual Contouring implement uniform spatial discretization strategies for surface extraction.
@@
We categorize existing methods into two distinct classes: Direct Extraction Methods, which utilize fixed information for isosurface extraction without additional input during reconstruction (including Marching Cubes, Dual Contouring, RFTS, and RFTA), and Iterative Refinement Methods, which progressively insert new points and adjust results during the extraction process (McGrids).
VI-B Comparison with Direct Extraction Methods
@@
Figure 6 illustrates comparative analyses at uniform grid resolutions of $30^{3}$ and $50^{3}$ over the $[-1,1]^{3}$ domain, where all algorithms using identical grid point SDF values and gradient information for low-genus model reconstruction. Marching Cubes and Dual Contouring utilize only local SDF information, resulting in poor detail preservation; RFTS cannot handle topological variations, restricting its application to simple models; and RFTA, despite improving upon its predecessor, fails to recover intricate surface details. Our method leverages projection points derived from gradient information with regular Delaunay-based extraction to achieve superior reconstruction quality. This is demonstrated in Figure 6 with the strawberry model and in Figure 2 with the golf ball model, where our approach successfully reconstructs surface cavities that RFTA fails to capture. Additionally, at $50^{3}$ resolution, RFTA’s tower reconstruction exhibits notable spherical artifacts.
@@
For high-genus models (Figure 7), we present two implementations: Our1, which uses standard $50^{3}$ grid inputs, and Our2, which employs an adaptive approach starting from an $8^{3}$ grid and expanding to $50^{3}$ total points (including both sample points and their projections) using the strategy detailed in Section V. While conventional methods (Marching Cubes, Dual Contouring, RFTS) fail to preserve intricate details such as chair backs, and RFTA shows only modest improvement, both of our implementations demonstrate better detail preservation. Our2, through its adaptive point sampling strategy, achieves particularly significant improvements in detail preservation.
@@
VI-C Comparison with Progressive Refinement Methods
@@
TABLE I: Quantitative comparison (CD $\cdot 10^{-5}$) for fixed grid inputs.
Resolution MC DC RFTS RFTA Ours1
303 267 252 244 11.5 2.68
403 140 114 258 8.27 1.30
603 68.4 60.9 236 4.93 0.48
TABLE II: Quantitative comparison on complex models.
[…]
MC 0.571 0.869 0.359 0.974 0.174 11.46
DC 0.864 0.794 0.352 0.945 0.167 61.63
RFTS 3.154 0.761 9.450 0.636 0.009 34.17
RFTA 3.050 0.836 1.485 0.729 0.007 $>$ 1h
McGrids 0.264 0.888 0.379 0.999 0.191 10.53
Ours2 0.216 0.897 0.348 0.999 0.187 10.15
VI-D Quantitative Evaluation
To comprehensively evaluate our method’s effectiveness, we conducted extensive quantitative comparisons with existing approaches. Our quantitative evaluation comprised two distinct experiments. First, we assessed extraction accuracy with fixed grid inputs—the core scenario addressed by RFTS/RFTA—at three different resolutions across 100 diverse models, with $L_{2}$ Chamfer Distance (CD) results presented in Table I. Second, we compared extraction quality on 50 complex shapes with intricate geometric features, where MC, DC, RFTS, and RFTA utilized 2003 uniform grids, while McGrids and our proposed method employed 60K adaptive samples, as shown in Table II. For all quantitative evaluations, we used 1 million sampled points to calculate the error metrics. The results demonstrate that our method achieves state-of-the-art extraction quality under fixed uniform sampling conditions while also producing superior surface reconstruction results in minimal computational time when utilizing adaptive SDF sampling strategies.
[…]
VI-E Computational Efficiency Analysis
Figure 9 presents a analysis of our algorithm’s computational performance, measured by evaluating execution time against the number of iteratively inserted sampling points on two topologically complex models. Our localized update strategy demonstrates significant efficiency advantages: the time required to insert each new point remains approximately constant regardless of the total number of points already incorporated into the regular Delaunay tetrahedralization. This locality property enables our method to achieve computational complexity that scales approximately linearly with point count ($O(n)$), avoiding the higher-order complexity typically associated with global mesh reconstruction approaches. The empirical measurements confirm that our incremental framework maintains consistent performance even as the sampling density increases, validating the effectiveness of our design decision to prioritize localized updates over global recomputation during the progressive refinement process.
VI-F Progressive Topological and Geometric Convergence Analysis
Figure 10 illustrates the progressive refinement of our algorithm on two representative models: one with high-genus topology and another with intricate surface details. Beginning with a uniform initialization of $15^{3}$ sampling points, the algorithm sequentially incorporates new sites until reaching the target of 50,000 sites. At each intermediate stage, highlighted sites indicate newly added positions in the transition to the subsequent refinement level. The visualization reveals two fundamental mechanisms of our adaptive site insertion strategy: (1) Topology correction—demonstrated in the high-genus model, where sites are strategically positioned to either establish critical connections or separate incorrectly merged regions; and (2) Geometric refinement—evidenced in the detailed model, where the progressive addition of sites systematically enhances surface feature resolution. These results confirm the efficacy of our quality-driven insertion approach in simultaneously improving topological accuracy and geometric fidelity through targeted, iterative refinement.
VI-G Ablation Study
To systematically evaluate the critical components of our method, we conducted comprehensive ablation studies on two fundamental aspects: (1) the contribution of projection points in the isosurface extraction process, and (2) the efficacy of our quality-driven adaptive point insertion strategy.
[…]
Figure 11: Comparison of results using (a) sampling points only versus (b) our complete method with both sampling points and their surface projections. Starting from a uniform $20^{3}$ grid and refining to 40,000 points, the projection-enhanced approach demonstrates superior topological preservation.
Impact of SDF Projection Points. Incorporating projection points into our power diagram computation significantly improves the algorithm’s capacity to accurately capture the underlying surface during isosurface extraction, particularly in regions with complex topological structures or fine geometric details. Figure 11 illustrates the comparative results using a uniform $15^{3}$ grid as initial input, iteratively refined to 40,000 sampling points with: (a) our baseline method using only sampling points without projections, and (b) our complete method integrating both sampling points and their corresponding surface projections.
The visual comparison reveals that the projection-enhanced approach successfully preserves topological integrity across the entire structure, while the baseline method exhibits substantial topological defects and missing features. These results demonstrate that surface projections provide crucial geometric cues that guide the power diagram construction toward more accurate surface representation. This guidance is especially valuable for reconstructing challenging features such as thin structures, sharp edges, and regions with high curvature variation that might otherwise be inadequately captured when relying solely on discrete SDF sampling points.
[…]
Effect of Adaptive Point Insertion Strategy on Topology Preservation. To quantify the effectiveness of our sequential point insertion strategy, we conducted a comparative analysis using varying initial uniform grid resolutions while maintaining a consistent total of 40,000 sampling points. As illustrated in Figure 12, reconstruction quality correlates directly with the proportion of adaptively inserted samples: when a higher percentage of points are strategically added through our quality-driven criteria, the model demonstrates significantly superior topological fidelity compared to reconstructions relying predominantly on uniform sampling. The uniform sampling approach produces reconstructions with notable topological artifacts and missing features, particularly in geometrically complex regions. These results empirically validate that our adaptive point insertion mechanism intelligently prioritizes topologically critical regions, optimizing computational resources to preserve essential geometric features with minimal sampling budget.
VI-H Neural SDF Extraction
Neural SDFs have gained widespread applications in computer graphics and vision. Although neural SDFs cannot strictly guarantee unit-length gradients, making projection points inaccurate, our method demonstrates robust extraction capabilities when applied to these representations. We validated our approach on two classic neural SDF frameworks: SIREN-based SDFs fitted from mesh data and VolSDF reconstructed from multi-view images. As shown in Figure 13, even with only 10,000 sampling points, our method successfully extracts complex structures from SIREN SDFs, demonstrating that our algorithm maintains its powerful topological preservation capabilities when applied to neural implicit representations.
Efficient extraction from neural SDFs requires effective parallelization to address computational overhead. We implemented a batch-based processing strategy where multiple points are inserted in sequence, followed by parallel evaluation of all affected regions before proceeding to the next insertion batch. This approach significantly reduces neural network query overhead by leveraging GPU parallelization, making our method computationally practical for neural SDF extraction tasks.
The effectiveness of our extraction strategy on neural SDFs stems from two complementary factors. First, our adaptive sampling mechanism prioritizes point placement near the zero-level set—precisely where neural SDFs tend to maintain higher accuracy. Second, the approximation errors in neural SDFs often exhibit a self-balancing property, where local inaccuracies in different regions tend to compensate for each other, resulting in globally coherent reconstructions when processed through our power diagram framework.
[…]
Figure 13: Neural SDF extraction in two representative scenarios: (1) SIREN-based SDFs fitted from mesh data and (2) VolSDF reconstructed from multi-view images.
VII Conclusion, Limitations and Future Work